# Understanding Parents’ Perceptions and Practices Toward Children’s Security and Privacy in Virtual Reality

### v1

@@ 4.1 Parents’ Perceptions of Children’s S&P Risks in VR (RQ1)
Due to our interview study’s small sample size in qualitative nature and our goal to highlight emerging insights as opposed to generalizability, we adopt a reporting method from a similar interview-based S&P study  [82] and use their frequency terminology throughout the remainder of the paper. Instead of using exact numbers of participants, we use “none” (0%), “a few” (0%-25%), “some” (25%-45%), “about half” (45%-55%), “most” (55%-75%), “almost all” (75%-100%), and “all” (100%), when appropriate.
In response to RQ1, our findings highlight: (1) Parents generally lacked S&P awareness in VR, largely due to the belief that VR is still in nascent development and therefore risks are a “future thing.” (2) S&P risks were regarded as potential risks, whereas the influence of VR on children’s physical health was defined as real-life risks by parents. (3) Among the identified S&P risks, parents generally had greater concerns in social interaction contexts than data surveillance by companies. Although they could still name a few risks from VR companies, they generally discarded such concerns in practice.
@@ 4.1.1 Non-S&P Concerns as Parents’ Primary or Only Concern for Children’s VR Usage
Defining eye-sight and physical injuries as real-life risks. When answering the question about general concerns for VR, most participants illustrated both risk-life risks they had perceived or experienced from their own children and potential risks they could imagine for other children or in future settings. For real-life risks, most parents mentioned concerns about the influence of VR on children’s eyesight. As P15 mentioned, “The only risk so far that we’ve had is my son needs glasses now, but we don’t know if there’s a direct correlation between the Oculus and his eyesight.
@@
Stranger danger in a child-adult co-existing ecosystem. Similar to what prior work has pointed out [4, 3, 40], the child-adult co-existence in VR was a major safety concern of parents. P1 mentioned:
@@
Some parents echoed P1, indicating their doubt about VR companies’ motive and capability to separate children from adult users. According to participants, the child-adult co-existence in the VR ecosystem could lead to many other potential problems, including predators and cyberbullying.
P6, an active social VR user, said, “I suppose there could be a way you could track where they (children) were actually if the person was a potential predator. I’m sure there’s a way to track the IP of a VR device.” P6 had heard of predators collecting IP addresses in social VR, and like some other parents also mentioned, they would be afraid of children being lured to meet online strangers in person, either voluntarily or involuntarily.
Besides predators, participants also mentioned that in an anonymous and immersive environment, cyberbullying could be common and more viscerally realistic. As P2 noted, “My largest concern would be, the potential for bullying within open environments where you can encounter cyberbullying, particularly with the potential lack of supervision.” Like P2, some participants were well aware of the anonymity afforded by avatar-mediated communication [83, 43, 84]. They perceived anonymous environments like VR to be a place where cyberbullying was more likely. Worse yet, as P11 stated, “[c]yberbullying becomes more visceral in VR (compared to text-based cyberbullying).”.
@@
Besides the under-regulated bad language, the age-mixed environment in VR also brought in the problem of inappropriate content, which has been documented in existing literature [3]. Participants, especially those with social VR experiences, said that the user-generated worlds were a big issue. Similar to the perception that VR companies did a bad job limiting the age group of users, participants also thought the mechanism of allowing users to upload their own worlds could potentially introduce children to content that was not designed for them.
Importantly, however, parents also noted that bad influences were not only from adults but also from other children who used VR without parents’ monitoring. P14, who did not have much social VR experience, said she could imagine “kids in VR might have families who have a different or opposite thought process from us.” Therefore, P14 was very concerned about the “bad influences” from other kids in VR.
@@ 4.1.2 S&P Concerns in Social Interaction Outweighed Data Surveillance
After asking participants about their general concerns, we specified our focus on children’s S&P in VR. About half of the participants said that they had never thought about the S&P issues in VR before. After being prompted, participants were able to name some S&P risks given their prior experience with other technologies. However, some participants would emphasize such concerns hardly occurred to them during children’s use of VR. In comparison, while some participants working in technology or S&P-related industries were more concerned about data surveillance from companies than self-disclosure in social interaction, most of the other participants were more concerned about children’s self-disclosure in VR.
Parents showed awareness of headsets’ and apps’ data collection risks but discarded the concerns in practice. For data S&P in VR, most parents mentioned their concerns for biometric data collection and analyses for user profiling and advertising purposes. Intriguingly, many children and family studies on S&P-related issues with technology have found that families easily neglect commercial data mining and surveillance in their S&P mental models [60, 16]. Our findings, however, show some progress that families have made in the awareness of surveillance capitalism, as our participants covered the four sensitive data types according to Dick [7], including observable data, observed data, computed data, and associated data in VR (see Section 2). For example, P20 said:
@@
For participants like P20, cutting-edge technology such as eye-tracking could collect physiological and behavioral signals, which could be more invasive compared to their smartphone activity. Echoing P20, P18 mentioned, “I would imagine there’s a lot more data that comes in when somebody’s interacting with VR, so I think that’s the biggest privacy risk is how much neural information you are giving up.” Besides “neural information”, participants were also concerned about their voice data being collected by companies, due to the expansion of AI that could be used to fabricate and manipulate voices.
Importantly, even though parents had an awareness of potential data collection by VR companies, they generally discarded such concerns in their parenting practices. As P1 highlighted,
@@
Although P1 acknowledged the risks of data collection by companies, he believed that those risks were still very abstract and theoretical to laypeople. He added, “A lot of people are doing a good job of presenting these pieces are scary but not doing a good job of telling me why they’re scary…But you know the risk of someone knowing your information.” We also noticed that a lot of what other participants described as VR data S&P risks testified to P1’s opinion. Participants knew that VR users gave out their biometric data but could not describe the consequences. However, they were able to accurately demonstrate example consequences (detailed below) of children’s self-disclosure in VR.
@@
In addition, some participants attributed children’s self-disclosure to the immersion and avatars. P20 said, “Feeling immersed in VR might make it a more likely scenario where children might disclose something to someone behind an avatar.” Compared to data collection by companies, participants could describe self-disclosure risks in VR in a much more vivid and specific way. They also considered such risks much more “direct”, despite acknowledging the subtle and potentially long-term risks of data surveillance in VR.
@@ 4.1.3 Reasons for Lack of Concerns: Being “Cutting-Edge” and “Harmless”
Parents lacked the motivation to fight a future thing. Some participants mentioned that VR was still a privilege in its infancy. For this reason, participants considered VR much less risky than some other social technologies that are being used by a majority of people. As P20 explained,
@@
Similarly, a large portion of participants mentioned that the risks they identified would not be concerning until some point in the future when VR technology becomes much more advanced and popular than its current state. Risks of VR are not “well-defined” at this point, as P1 said:
@@
P1 showed a state of mind inconsistent with his VR parenting practice, demonstrating the existence of privacy paradox [85, 86]. Privacy paradox refers to the discrepancies between users’ expressed privacy concerns and their actual behavior. In our context, even though P1 knew the risks existed, he still had little motivation to mitigate them in practice.
As a result, parents usually focused more on their children’s use of daily technologies like social media, neglecting the harms of “future technology” – VR.
Trust in children’s maturity in technology use. In addition to defining VR risks as a future thing, some participants, most of whom were parents of teenagers (aged 13-17), said that they were not concerned because their children were tech-savvy and mature in their decision-making. As P14 said, “I think at this point for him. He’s kind of on his path to becoming an adult. So I’ve given him advice, and I’m kind of trusting that he’s gonna follow it. He’s fairly mature for 16.” Similarly, other parents of children at a similar age echoed P14 and believed their children’s familiarity with VR to have surpassed theirs.
Some participants who were experienced in VR said that they suspected most other parents would have bought VR just for babysitting their children without having any knowledge of VR. As P1 said, “I don’t think parents understand that there’s a shift there. It goes from being something that’s happening on a screen in front of you to something that is happening to you.” P1 assumed that many parents would consider VR the same as many other technologies like phones, TVs, and tablets.
Indeed, when other participants discussed some risks of VR, they would reference other daily technologies, like what P14 argued, “the risks are pretty much the same in any type of technology.
Some participants not only neglected the nuances between different technologies, but they also regarded VR as a kind of video game that was just the same as being played on a PC. Interestingly, participants were generally much less concerned about their S&P when in gaming contexts. As P16 explained why she did not worry about her son’s privacy in VR, “I am not concerned because they’re just in this game mode. There’s nothing personal involved.” For some participants like P16, gaming sessions were mostly about play and involved very few privacy risks, hence lacking concerns.
Such an imbalanced understanding of technology between parents and children has also been widely documented in existing literature [23, 24], impeding parents’ ability to engage in teens’ technology use.
@@ 4.2 Parents’ S&P Risk Mitigation Strategies (RQ2)
Following Wisniewski et al.’s TOSS framework [22], our findings highlight: (1) Parents primarily relied on active mediation, rules, and restrictions in usage. (2) Parents generally considered parental controls in VR important but faced significant technical constraints. (3) In mitigating S&P risks, parents had to navigate complex tradeoffs between autonomy and safety.
@@ 4.2.1 Active Strategies as the Primary Strategy: Proactively Mitigating Risks via Conversations and Rules
We asked participants what steps they had taken to address the concerns they had, and almost all parents mentioned proactively educating children through conversations and setting rules. However, most of them also emphasized that these conversations and rules were not exclusive to their VR usage, hence neglecting their ineffectiveness in certain contexts.
@@
Participants like P14 who said they had strict rules and open conversations with their children all claimed their children followed their guidance well. However, P14 also emphasized that their rules were “pretty much the same across any kind of electronic communication that can go outwards.
@@
According to P7, although parents might have had sufficient privacy education in general digital communications, the proximity between VR and real-life interactions makes it more reasonable to set physical references in their VR privacy education, as opposed to other digital communications in general.
Managing children’s app access and S&P information flows in VR through shared and associated accounts. Unlike social media usage where children have separate accounts from parents [87], most participants shared VR headsets and accounts with their children. While this practice was not out of mitigation considerations, some participants indicated that sharing headsets and accounts helped them better manage children’s VR activity, including their access to certain apps. For example, P14 mentioned:
@@
In some cases, participants made sure to be present when their children tried to create accounts, in case of providing excessive information for platforms. As P10 noted, “I’m concerned about the phone number and other family information going out. That’s why I keep my kids away from creating accounts on their own without having us know.
@@ 4.2.2 Challenges in Passive Strategies: Technical Issues & Parents’ Frustrations
Maintaining ambient awareness  [^1] Referencing [16], we use this term to indicate parents’ casual monitoring when they are nearby their children when using technology. as the most common monitoring strategy. Almost all participants said they would occasionally check in with children when playing VR, making sure to be able to hear the conversations. In addition, for physical safety considerations, most participants would ask their children to play in open areas like the living room. For example, P11 said:
@@
However, unless casting VR on a screen, parents would not be able to see what their children do. While paying attention to the video sounds is also a more common option for parents who monitor children using 2D media [16], VR was much less flexible for our participants who would like to visually check out their children. This constraint influenced parents’ practices and attitudes toward different parental controls in VR, as we discuss below.
Parental controls in VR were important but posed technical constraints. In the last section of our interviews, we asked participants to first indicate if they know the feature, and then rate their anticipated effectiveness of each parental control that we found in mainstream VR products.
@@
Some participants like P20 also said that they had never tried setting up parental controls, but given their experience with other technology mediation, they assumed they would be able to set them up.
@@
Besides headset mirroring, blocking apps and content were also highly rated, due to their preventive nature favored by our participants. However, a few participants also pointed out the limitation. As P17 mentioned, “Blocking is not going to be effective. There’s always going to be something that you don’t know about.
To our surprise, while social communication was the most concerning privacy risk to participants, restricting it was not the most effective feature for participants. As P18 explained, “I mean, it’s a question that, if you want your child to be able to communicate socially, it’ll be a kind of an all-or-nothing thing.” While restricting social communication could prevent children from interacting with strangers, it was also perceived as too limiting and could not “help them learn to do things in the correct way.” (P18). Indeed, passive mediation means stopping children from doing certain things while hindering their chances of learning in practice. Navigating such tradeoffs between children’s S&P and their autonomy was a hard process for our participants (more detailed in 4.2.3).
@@
Participants who preferred to trade autonomy for children’s S&P in VR considered the latter more important mainly for two reasons. First, the current VR ecosystem is largely designed for adults and children are more likely to get into inappropriate situations at this point than later when VR gets more mature. Hence, it is still too early to give children too much autonomy. Second, VR is still a very minor part of most children’s lives, and therefore they have plenty of opportunities to maintain their social autonomy elsewhere, e.g., physical interaction contexts, smartphone usage, etc. As P1 noted:
@@
On the other hand, participants who advocated for children’s autonomy even at the cost of safety and S&P risks believed it would be more important for children to gradually learn to tackle harmful situations with parents’ guidance. Using passive strategies would hinder children’s chances of learning in the real world where parents cannot always keep them company.
Consequently, some participants advocating for the second stance claimed that parents could only help them understand and prepare for potential risks beforehand, whereas when real risks took place, they had to count on themselves with more effective moderation or defensive mechanisms to be in place – a gap that VR companies have the responsibilities to fill (detailed in 4.3). Importantly, most participants also emphasized that the balance between autonomy and children’s S&P and safety in VR would be a dynamic process where the significance of one and the other might be constantly changing in the long run.

### v3

@@ 4.1 Parents’ Perceptions of Children’s S&P Risks in VR (RQ1)
Due to our interview study’s small sample size in qualitative nature and our goal to highlight emerging insights as opposed to generalizability, we adopt a reporting method from a similar interview-based S&P study  [89] and use their frequency terminology throughout the remainder of the paper. Instead of using exact numbers of participants, we use “none” (0%), “a few” (1%-24%), “some” (25%-44%), “about half” (45%-54%), “most” (55%-74%), “almost all” (75%-99%), and “all” (100%), when appropriate.
In response to RQ1, our findings highlight: (1) Parents generally lack S&P awareness in VR, largely due to the belief that VR is still in nascent development and therefore risks are a “future thing.” (2) S&P risks are regarded as potential risks, whereas the influence of VR on children’s physical health is defined as real-life risks by parents. (3) Among the identified S&P risks, parents generally have greater concerns in social interaction contexts than data surveillance by companies. Although they can name a few risks from VR companies, they generally discard such concerns in practice.
@@ 4.1.1 Non-S&P Concerns as Parents’ Primary or Only Concern for Children’s VR Usage
Defining eye-sight and physical injuries as real-life risks. When answering the question about general concerns for VR, most participants illustrated both risk-life risks they had perceived or experienced from their own children and potential risks they could imagine for other children or in future settings. For real-life risks, most participants mentioned concerns about the influence of VR on children’s eyesight. As P15 mentioned, “The only risk so far that we’ve had is my son needs glasses now, but we don’t know if there’s a direct correlation between the Oculus and his eyesight.
@@
Stranger danger in a child-adult co-existing ecosystem. Similar to what prior work has pointed out [4, 3, 42], child-adult co-existence in VR was a major safety concern. P1 mentioned:
@@
Some participants echoed P1, indicating their doubt about VR companies’ motive and capability to separate children from adult users. According to participants, the child-adult co-existence in the VR ecosystem could lead to many other potential problems, including predators and cyberbullying.
P6, an active social VR user, said, “I suppose there could be a way you could track where they (children) were actually if the person was a potential predator. I’m sure there’s a way to track the IP of a VR device.” P6 had heard of predators collecting IP addresses in social VR, and like some other participants also mentioned, they would be afraid of children being lured to meet online strangers in person, either voluntarily or involuntarily.
Besides predators, participants also mentioned that in an anonymous and immersive environment, cyberbullying could be common and more realistic. As P2 noted, “My largest concern would be, the potential for bullying within open environments where you can encounter cyberbullying, particularly with the potential lack of supervision.” Like P2, some participants were well aware of the anonymity afforded by avatar-mediated communication [90, 45, 91]. They perceived anonymous environments like VR to be a place where cyberbullying was more likely. Worse yet, as P11 stated, “[c]yberbullying becomes more visceral in VR (compared to text-based cyberbullying).”.
@@
Besides the under-regulated bad language, the age-mixed environment in VR also brought in the problem of inappropriate content, which has been documented in existing literature [3]. Participants, especially those with social VR experience, said that the user-generated worlds were a big issue. Similar to the perception that VR companies did a bad job limiting the age group of users, participants also thought the mechanism of allowing users to upload their own worlds could potentially introduce children to content that was not designed for them.
Importantly, however, participants also noted that bad influences were not only from adults but also from other children who used VR without parents’ monitoring. P14, who did not have much social VR experience, said she could imagine “kids in VR might have families who have a different or opposite thought process from us.” Therefore, P14 was very concerned about the “bad influences” from other kids in VR.
@@ 4.1.2 S&P Concerns in Social Interaction Outweighed Data Surveillance
After asking participants about their general concerns, we specified our focus on children’s S&P in VR. About half of the participants said that they had never thought about the S&P issues in VR before. After being prompted, participants were able to name some S&P risks given their prior experience with other technologies. However, some participants would emphasize such concerns hardly occurred to them during children’s VR usage. In comparison, while some participants working in technology or S&P-related industries were more concerned about data surveillance from companies than self-disclosure in social interaction, most of the other participants were more concerned about children’s self-disclosure in VR. Mapping this finding to the threat model suggested by Garrido et al. [40], we suggest that parents with more S&P experience are more concerned about hardware and server adversaries. Still, most parents are much more concerned about user adversaries.
Parents showed awareness of headsets’ and apps’ data collection risks but discarded the concerns in practice. For data S&P in VR, most parents mentioned their concerns for biometric data collection and analyses for user profiling and advertising purposes. Intriguingly, many children and family studies on S&P-related issues with technology have found that families easily neglect commercial data mining and surveillance in their S&P mental models [62, 16]. Our findings, however, show some progress that families have made in developing awareness of surveillance capitalism, as our participants covered the four sensitive data types according to Dick [7], including observable data, observed data, computed data, and associated data in VR (see Section 2). For example, P20 said:
@@
For participants like P20, cutting-edge technology such as eye-tracking could collect physiological and behavioral signals, which could be more invasive compared to their smartphone activity. Echoing P20, P18 mentioned, “I would imagine there’s a lot more data that comes in when somebody’s interacting with VR, so I think the biggest privacy risk is how much neural information you are giving up.” Besides “neural information,” participants were also concerned about their voice data being collected by companies, due to the expansion of AI that could be used to fabricate and manipulate voices.
Importantly, even though participants were aware of the potential data collection and inference by VR companies, they generally discarded such concerns in their parenting practices. As P1 highlighted,
@@
Although P1 acknowledged the risks of data collection and inference by companies, he believed that those risks were still very abstract and theoretical to laypeople. He added, “A lot of people are doing a good job of presenting these pieces are scary but not doing a good job of telling me why they’re scary…But you know the risk of someone knowing your information.” We also noticed that a lot of what other participants described as VR data S&P risks testified to P1’s opinion. Participants knew that VR users gave out their biometric data but could not describe the consequences. However, they were able to accurately demonstrate example consequences (as detailed below) of children’s self-disclosure in VR.
@@
In addition, some participants attributed children’s self-disclosure to the immersion and avatars. P20 said, “Feeling immersed in VR might make it a more likely scenario where children might disclose something to someone behind an avatar.” Compared to data collection and inference by companies, participants could describe self-disclosure risks in VR in a much more vivid and specific way. They also considered such risks much more “direct”, despite acknowledging the subtle and potentially long-term risks of data surveillance in VR.
@@ 4.1.3 Reasons for Lack of Concerns: Being “Cutting-Edge” and “Harmless”
Parents lacked the motivation to fight a future thing. Some participants mentioned that VR was still a privilege in its infancy. For this reason, participants considered VR much less risky than some other technologies that are being used by a majority of people. As P20 explained,
@@
Similarly, a large portion of participants mentioned that the risks they identified would not be concerning until some point in the future when VR technology becomes much more advanced and popular than its current state. VR risks are not “well-defined” at this point, as P1 said:
@@
P1 showed a state of mind that is inconsistent with his VR parenting practice, demonstrating privacy paradox [92, 93]. Privacy paradox refers to the discrepancies between users’ expressed privacy concerns and their actual behavior. In our context, even though P1 knew the risks existed, he still had little motivation to mitigate them in practice.
As a result, participants usually focused more on their children’s use of other technologies like social media, neglecting the harms of VR as a “future technology.”
Trust in children’s maturity in technology use. In addition to defining VR risks as a future thing, some participants, especially those who were parents of teenagers (aged 13-17), said that they were not concerned because their children were tech-savvy and mature in their S&P decision-making. As P14 said, “I think at this point for him. He’s kind of on his path to becoming an adult. So I’ve given him advice, and I’m kind of trusting that he’s gonna follow it. He’s fairly mature for 16.” Similarly, other parents of teenagers echoed P14 and believed their children’s familiarity with VR to have surpassed theirs.
Some participants who were experienced in VR said that they suspected most other parents would have bought VR just to babysit their children without having any knowledge of VR. As P1 said, “I don’t think parents understand that there’s a shift there. It goes from being something that’s happening on a screen in front of you to something that is happening to you.” P1 assumed that many parents would consider VR the same as many other technologies like phones, TVs, and tablets.
Indeed, when other participants discussed some risks of VR, they would reference more common technologies, like what P14 (a parent of teenagers) argued, “the risks are pretty much the same in any type of technology.
Some participants not only neglected the nuances between different technologies, but they also regarded VR as a kind of video game that was just the same as being played on a PC. Interestingly, participants were generally much less concerned about their S&P when in gaming contexts. As P16, a parent of a teenager nearing adulthood, explained why she did not worry about her son’s privacy in VR, “I am not concerned because they’re just in this game mode. There’s nothing personal involved.” For some participants like P16, gaming sessions were mostly about play and involved very few privacy risks.
In contrast, parents of children at younger ages (below 13) in our study, such as P2, P7, and P15, showed more vigilance toward VR risks. Most of the parents discussed that they changed or planned to change their strategies as children grew older. Consequently, some parents of older children, especially those nearing adulthood, might be more neglectful of VR risks compared to those of younger children. In general, parents of older children demonstrated an even more imbalanced understanding of VR compared to their children.
Such information asymmetry about technology between parents and children has also been widely documented in existing literature [23, 24], impeding parents’ ability to engage in teens’ technology use.
@@ 4.2 Parents’ S&P Risk Mitigation Strategies (RQ2)
Following Wisniewski et al.’s TOSS framework [22], our findings highlight: (1) Parents primarily rely on active mediation, rules, and restrictions in usage. (2) Parents generally consider parental controls in VR important but face significant technical constraints. (3) In mitigating S&P risks, parents have to navigate complex tradeoffs between autonomy, safety, and S&P.
@@ 4.2.1 Active Strategies as the Primary Strategy: Proactively Mitigating Risks via Conversations and Rules
We asked participants what steps they had taken to address the concerns they had, and almost all participants mentioned proactively educating children through conversations and setting rules. However, most of them also emphasized that these conversations and rules were not exclusive to their VR usage. Hence, the ineffectiveness of such strategies in certain VR contexts might have been neglected.
@@
Participants like P14 who said they had strict rules and open conversations with their children all claimed that their children followed their guidance well. However, P14 also emphasized that their rules were “pretty much the same across any kind of electronic communication that can go outwards.
@@
According to P7, although parents might have had sufficient privacy education in general digital communication, the proximity between VR and real-life interactions makes it more reasonable to set physical references in their VR privacy education, as opposed to sharing the same set of rules across all digital communication in general.
Managing children’s app access and S&P information flows in VR through shared and associated accounts. Unlike social media usage where children usually have separate accounts from parents [94], most participants shared the only set of VR headsets and accounts with their children. While this practice was not out of mitigation considerations, some participants indicated that sharing headsets and accounts helped them better manage children’s VR activity, including their access to certain apps. For example, P14 mentioned:
@@
In some cases, participants made sure to be present when their children tried to create accounts in case they provided excessive information for platforms. As P10 noted, “I’m concerned about the phone number and other family information going out. That’s why I keep my kids away from creating accounts on their own without having us know.
@@
Notably, this approach was more discussed by parents of children aged below 13. For parents of children aged above 13, almost all of them did not mention managing or sharing accounts with their children. A few of them mentioned that they would only occasionally check their activity (e.g., interactions with other users) without overly intruding on children’s privacy. In addition, the only two participants whose children had their own headsets were aged above 13, further indicating that parents of teenagers may have fewer restrictions and monitoring over their VR usage. This shows a gap where parents of children aged above 13 have limited concerns and mediation toward their children’s use of VR products that primarily encourage users aged above 13 [86, 87, 88].
@@ 4.2.2 Challenges in Passive Strategies: Technical Issues & Parents’ Frustrations
Maintaining ambient awareness  [^1] Referencing [16], we use this term to indicate parents’ casual monitoring when they are nearby their children during technology use. as the most common monitoring strategy. Almost all participants said they would occasionally check in with children when playing VR, making sure to be able to hear the conversations. In addition, for physical safety considerations, most participants would ask their children to play in open areas like the living room. For example, P11 said:
@@
However, unless VR is cast on a screen, participants would not be able to see what their children do. While paying attention to the sounds is also a more common option for parents who monitor children using 2D media [16], VR was much less flexible for our participants who would like to visually check out their children. This constraint influenced parents’ practices and attitudes toward different parental controls in VR, as we discuss below.
Parental controls in VR were important but posed technical constraints. In the last section of our interviews, we asked participants to first indicate if they knew the feature, and then rate their anticipated effectiveness of each parental control that we found in mainstream VR products.
@@
Some participants like P20 also said that they had never tried setting up parental controls, but given their experience in other technology mediation, they assumed they would be able to set them up.
@@
Besides headset mirroring, blocking apps and blocking content were also highly rated, due to their preventive nature favored by some participants. However, a few participants also pointed out the limitation. As P17 mentioned, “Blocking is not going to be effective. There’s always going to be something that you don’t know about.
To our surprise, while social communication was the most concerning privacy risk to participants, restricting it was not the most effective feature for participants. As P18 explained, “I mean, it’s a question that, if you want your child to be able to communicate socially, it’ll be a kind of an all-or-nothing thing.” While restricting social communication could prevent children from interacting with strangers, it was also perceived as too limiting and could not “help them learn to do things in the correct way.” (P18). Indeed, passive mediation means stopping children from doing certain things and hindering their chances of learning in practice. Navigating such tradeoffs between children’s S&P and their autonomy was a hard process for our participants (as detailed in Section 4.2.3). In general, we observed that parents of children aged above 13 were more against the idea of using passive features that could hinder children’s autonomy. Hence, some parental controls were much less important to parents of children aged above 13. On the other hand, parents of younger children favored these features for their effectiveness in protecting children from harms.
@@
Participants who preferred to trade autonomy for children’s S&P in VR considered the latter more important mainly for two reasons. First, the current VR ecosystem is largely designed for adults, and children are more likely to get into inappropriate situations at this point than later when VR becomes more mature. Hence, it is still too early to give children too much autonomy. Second, VR is still a very minor part of most children’s lives, and therefore they have plenty of opportunities to maintain their social autonomy elsewhere, e.g., physical interaction contexts, smartphone usage, etc. As P1 noted:
@@
On the other hand, participants who advocated for children’s autonomy even at the cost of safety and S&P believed it would be more important for children to gradually learn to tackle harmful situations with parents’ guidance. Using passive strategies would hinder children’s chances of learning in the real world where parents cannot always keep them company.
Consequently, some participants advocating for the second stance claimed that parents could only help them understand and prepare for potential risks beforehand, whereas when real risks took place, they had to count on themselves with more effective moderation or defensive mechanisms to be in place – a gap that VR companies have the responsibilities to fill (as detailed in Section 4.3). Importantly, most participants also emphasized that the balance between autonomy and children’s S&P and safety in VR would be a dynamic process where the significance of one and the other might be constantly changing in the long run of children’s development. Specifically, almost all of the participants who valued S&P over autonomy had children aged below 13, whereas almost all of the participants whose children were aged above 13 valued autonomy more. This shows the role of children’s age in impacting parents’ decision-making when facing the autonomy-S&P tradeoffs.